Suppose the marginal cost of writing a contract of length L is MC(L) = 10 + 2L. Find the optimal contract length when the marginal benefit of writing a contract is: Instruction: Enter your responses rounded to two decimal places. a. MB(L) = 100. b. MB(L) = 150. c. What happens to the optimal contract length when the marginal benefit of writing a contract increases? It stays the same. It increases. It decreases.
Answer #1

At optimal point, MC(L) = MB(L)

(a) MB(L) = MC(L)

100 = 10 + 2L

2L = 100 -10

2L = 90

L = 90/2

L = 45. (Optimal Contract Length)

(b) MB(L) = MC(L)

150 = 10 + 2L

2L = 150 -10

2L = 140

L = 140 / 2

L = 70 (Optimal contract length)

(c) Optimal contract length rises from 45 to 70 when the the marginal benefit of writing a contract increases.

Answer: It increases.